\begin{thebibliography}{}\label{sec:TeXbooks}
\bibitem[*]{email}{Electronic Address: lake@astro.queensu.ca}
\bibitem{old} Two classic examples are the discussions in the text by Tolman (R.C. Tolman, {\it
Relativity, Thermodynamics and Cosmology}, (Oxford Univ. Press,
London, 1934)) and in the review by Robertson (H. P. Robertson,
Rev. Mod. Phys., \textbf{5}, 62 (1933)).
\bibitem{current} For example, a search at arXiv.org of ``bouncing
cosmologies" returns over 100 abstracts at the time of writing.
\bibitem{Steinhardt}See, for example, P. J. Steinhardt and N.
Turok, ``The Cyclic Universe: An Informal Introduction"
(astro-ph/0204479).
\bibitem{visser1}See C. Molina-Paris and M. Visser, Phys. Lett. B, \textbf{455},
90 (1999) (gr-qc/9810023) and, for a less general analysis,  C.P.
Burgess, F. Quevedo, R. Rabadan, G. Tasinato and I. Zavala,  JCAP
0402 (2004) 008. (hep-th/0310122)
\bibitem{Riess}A. G. Riess \textit{et al.}, ``Type Ia Supernova
Discoveries at $z>1$ From the Hubble Space Telescope: Evidence for
Past Deceleration and Constraints on Dark Energy Evolution", Ap.
J. \textbf{607}, 665 (2004). (astro-ph/0402512)
\bibitem{Riess98}A. G. Riess \textit{et al.}, A. J., \textbf{116}
1009 (1998). (astro-ph/9805201)
\bibitem{Perlmutter} S. Perlmutter \textit{et al.}, Ap. J., \textbf{517},
565 (1999). (astro-ph/9812133)
\bibitem{wmap}D. N. Spergel \textit{et al.}, ``First Year Wilkinson
Microwave Anisotropy Probe (WMAP) Observations: Determination of
Cosmological Parameters", Ap. J. Suppl. \textbf{148}, 175 (2003)
(astro-ph/0302209)
\bibitem{lake1} See, for example, K. Lake, ``Cosmological observations and their representation in the $\Omega_{\Lambda}-\Omega_{M}$ plane"
(2004). (astro-ph/0404319)
\bibitem{rw}The uniqueness of the FLRW metric can be derrived
without any reference to field equations and in particular without
referenece to any background mathematical fluid. In this regard it
is perhaps useful to view the metric (\ref{rw}) as derived from
exact spherical symmetry about every point. See A. G. Walker, J.
London. Math. Soc., \textbf{19}, 219 (1944).
\bibitem{conventions} We use a signature of
+2 and start with geometrical units. Latin indicies range from 1
to 3 and Greek indicies from 1 to 4.
\bibitem{invariants}See, for example, K. Santosuosso, D. Pollney,
N. Pelavas, P. Musgrave and K. Lake, Computer Physics
Communications \textbf{115}, 381 (1998). (gr-qc/9809012)
\bibitem{kretschmann}In
particular, there is no need to look at the traditional
``Kretschmann" invariant $K=R_{\alpha \beta \gamma
\delta}R^{\alpha \beta \gamma \delta}$ ($R_{\alpha \beta \gamma
\delta}$ the Riemann tensor).
\bibitem{lovelock}D. Lovelock, J. Math. Phys. \textbf{13}, 874 (1972).
\bibitem{lake2} For use of these distinctions in a different context see
K. Lake, ``Sudden future singularities in FLRW cosmologies",
Classical and Quantum Gravity (in press) (gr-qc/0407107)
\bibitem{ellis} See, for example, \textit{Dynamical Systems in
Cosmology} Edited by J. Wainwright and G. F. R. Ellis (Cambridge
University Press, Cambridge, 1997).
\bibitem{unique} For further discussion see H. Stephani, D. Kramer, M.
MacCallum, C. Hoenselaers and E. Herlt, \textit{Exact Solutions of
Einstein's Field Equations} (Cambridge University Press,
Cambridge, 2003) and A. Krasi\'{n}ski, \textit{Inhomogeneous
Cosmological Models} (Cambridge University Press, Cambridge,
1997).
\bibitem{friedmann} Friedmann's papers, with expert and referenced editorial notes,
have been reprinted in English translation in \textit{Gen. Rel.
Grav.} \textbf{31}, 1985 (1999).
\bibitem{weinberg} See, for example, S. Weinberg astro-ph/9610044
and citations thereto and T. Padmanabhan, Phys.Rept.
\textbf{380},235 (2003). (hep-th/0212290).
\bibitem{he}S. W. Hawking and G. F. R. Ellis, \textit{The Large Scale Structure of Space-Time}
(Cambridge University Press, Cambridge, 1973)
\bibitem{visser} M. Visser, \textit{Lorentzian Wormholes} (Springer-Verlag,
New York, 1996)
\bibitem{poisson} E. Poisson, \textit{A
Relativist's Toolkit: The Mathematics of Black-hole Mechanics}
(Cambridge University Press, Cambridge, 2004).
\bibitem{clarification} These conditions are usually stated as
inequatities that hold for any vector field of a stated type. By
this we understand that the inequalities must hold over the set of
all possible vector fields of a stated type. The inequalities are
in general not appropriate for a subset as the associated physics
can be entirely regular for the subset under inequalities that
differ from the stated conditions.
\bibitem{note1}Note that the existence of this interval (without
$\mathcal{C}$) is required if the spacetime is to satisfy the null
energy condition.
\bibitem{maxima} At regular maxima consider $\ddot{a}<0$ in the
foregoing and it follows that only for $k=1$ can $\mathcal{C}$ be
$>0$ in the sufficient condition for $\hbox{EC}$.
\end{thebibliography}

\documentclass[aps,floatfix,twocolumn]{revtex4}
\usepackage{graphics,epsfig}
\usepackage{graphicx}

\begin{document}

\title{Energy conditions and a bounce in FLRW cosmologies}
\author{Benjamin K. Tippett and Kayll Lake \cite{email}}
\affiliation{Department of Physics, Queen's University, Kingston,
Ontario, Canada, K7L 3N6 }
\date{\today}

\begin{abstract}
Exploiting the existence of two ``cosmological" constants, one
associated with the classical Lovelock theorem and one with the
vacuum energy density, we argue, in a model independent way, that
in spatially closed FLRW cosmologies with a positive definite
effective cosmological constant there exists a range in this
constant that serves as a sufficient condition for the
satisfaction of the null, weak, strong  and dominant energy
conditions at a bounce. The application of energy conditions is
not unambiguous and we show how the bounce can be considered
classically and how, we believe more reasonably, it can be
considered a matter of quantum cosmology.
\end{abstract}

\maketitle

\textit{Introduction.} The idea of a bounce in
Friedmann-Lema\^{i}tre-Robertson-Walker (FLRW) cosmologies is an
idea that is as old as spacetime cosmology itself \cite{old} but
also one with current appeal \cite{current} motivated primarily by
M-theory, branes and extra-dimensions \cite{Steinhardt}. A bounce,
as opposed to a big bang, obviously meets with many problems when
trying to reproduce the universe we observe. Moreover, there is a
widely held belief that a bounce must violate classical energy
conditions \cite{visser1}. However, the application of classical
energy conditions is not unambiguous and here we exploit the
existence of two ``cosmological" constants, one associated with
the classical Lovelock theorem and one with the vacuum energy
density, and argue, in a model independent way, that in spatially
closed FLRW cosmologies with a positive definite effective
cosmological constant (as explained precisely below) there exists
a range in this constant that serves as a sufficient condition for
the satisfaction of the null, weak, strong  and dominant energy
conditions at a bounce. Further, depending on how the energy
conditions are applied, we show how a bounce might be considered
classically and how, we believe more correctly, it must be
considered a matter of quantum cosmology. In this analysis we
offer no explanation as to why there should be a bounce but rather
only examine the consequences of a bounce as regards the
application of energy conditions.

 The suggestion that we live in a
spatially closed universe with a positive definite observed
cosmological constant is in accord with the recent Hubble Space
Telescope observations of Riess \textit{et al.} \cite{Riess},
observations that not only confirm earlier reports that we live in
an accelerating universe \cite{Riess98} \cite{Perlmutter}, but
also sample the transition from deceleration to acceleration. In
contrast, observations with the Wilkinson Microwave Anisotropy
Probe \cite{wmap} argue strongly that the FLRW cosmology is
spatially flat  \cite{lake1}. In the spatially flat case we argue
that a bounce necessarily violates energy conditions no matter how
they are applied. Because our application of the energy conditions
differs in a significant way from previous works, we start our
discussion from first principles.

\bigskip
\maketitle \textit{Background Geometry.} The Cosmological
Principle is the working hypothesis that the Universe is described
by a spacetime that is spatially isotropic globally. It follows
from the theorems of Robertson and of Walker \cite{rw} that the
spacetime is represented by by the
Friedmann-Lema\^{i}tre-Robertson-Walker (FLRW) metric, which, up
to a coordinate transformation, is given by \cite{conventions}
\begin{equation}
ds^2=a^2d\bar{s}_{\Gamma}^2-dt^2 \label{rw}
\end{equation}
where $a=a(t)$, normally $\in C^2$, and $d\bar{s}_{\Gamma}^2$ is
the metric of a three space of constant curvature. In  $\Gamma$ we
scale the coordinates so that $k \equiv R_{a}^{b}R_{b}^{a}/2
(R_{a}^{a})^2 = \pm1,0$ where $R_{a}^{b}$ is the Ricci tensor of
$\Gamma$.

\bigskip

The metric (\ref{rw}) has only two independent (scalar) invariants
derivable from the Riemann tensor without differentiation
\cite{invariants}. These are usually taken to be the Ricci scalar
$R$ ($\equiv R_{\alpha}^{\alpha}$, $R_{\alpha}^{\beta}$ the Ricci
tensor of the spacetime) and first Ricci invariant $R1$ ($\equiv
S_{\alpha}^{\beta}S_{\beta}^{\alpha}/4 \equiv S^2$,
$S_{\alpha}^{\beta}$ the trace-free Ricci tensor $\equiv
R_{\alpha}^{\beta}-\delta_{\alpha}^{\beta}R/4$). Without any loss
in generality we can also take the two independent invariants to
be $\rho$ and $p$ defined as follows:
\begin{equation}
8 \pi \rho \equiv R/4+sign(S)\sqrt{3 S^2} -\mathcal{C} \label{rho}
\end{equation}
and
\begin{equation}
8 \pi p \equiv -R/4+sign(S)\sqrt{S^2/3} +\mathcal{C},  \label{p}
\end{equation}
where $\mathcal{C}$ is a constant. In the coordinates of
(\ref{rw}) these evaluate to
\begin{equation}
8 \pi \rho = \frac{3}{a^2}(k+\dot{a}^2) -\mathcal{C},
\label{rhoeval}
\end{equation}
and
\begin{equation}
8 \pi p = -\frac{1}{a^2}(k+\dot{a}^2+2a\ddot{a}) +\mathcal{C},
\label{peval}
\end{equation}
respectively where $^{.}\equiv d/d t$. Since $\rho$ and $p$ are
invariants, they distinguish scalar polynomial singularities of
the metric (\ref{rw}) directly \cite{kretschmann}.

\bigskip
\textit{Physical Interpretation.} A remarkable feature of four
dimensions is the fact that the only divergence free two index
tensor $A^{\alpha}_{\beta}$ derivable from the metric tensor and
its first two derivatives is (up to a disposable multiplicative
constant) \cite{lovelock}
\begin{equation}
A^{\alpha}_{\beta}=G_{\alpha}^{\beta}+ \lambda_{c}
\delta_{\alpha}^{\beta}, \label{lovelock}
\end{equation}
where $G_{\alpha}^{\beta}$ is the Einstein tensor ($\equiv
R_{\alpha}^{\beta}-\frac{1}{2}R \delta_{\alpha}^{\beta}$) and
$\lambda_{c} $ is a constant. We write Einstein's equations in the
form
\begin{equation}
G_{\alpha}^{\beta}+\lambda_{c} \delta_{\alpha}^{\beta}=8 \pi
_{M}T_{\alpha}^{\beta} + 8 \pi _{V}T_{\alpha}^{\beta},
\label{einstein}
\end{equation}
where $_{M}T_{\alpha}^{\beta}$ signifies the matter contribution
to (\ref{rw}) and $_{V}T_{\alpha}^{\beta}$ the vacuum
contribution. Here we simply assume that the latter is of the
usual form $8 \pi_{V}T_{\alpha}^{\beta} \equiv \lambda_{v}
\delta_{\alpha}^{\beta}$ where $\lambda_{v}$ a constant. Central
to our considerations here is the treatment of the constant
$\lambda_{c}$ which, it is fair to say, is usually ignored.

\bigskip

It is useful at this point, prior to the application of energy
conditions, to review three distinct ways of proceeding with
(\ref{einstein}) \cite{lake2}. Because of the form of (\ref{rw})
we can always consider the source a perfect mathematical fluid
(including, for example, scalar fields \cite{ellis}) with
associated streamlines comoving with $\Gamma$ \cite{unique}. The
flow (congruence of unit timelike vectors $u^{\alpha}$) and normal
field $n^{\alpha}$ (in the tangent space of the associated
Lorentzian two-space) are uniquely determined. The energy density
and isotropic pressure (including bulk viscosity) are now defined
by
\begin{equation}
\rho \; \equiv \; T_{\alpha}^{\beta}
u^{\alpha}u_{\beta}\label{rhodef}
\end{equation}
and
\begin{equation}
p \; \equiv \; T_{\alpha}^{\beta} n^{\alpha}n_{\beta}\label{pdef}
\end{equation}
respectively where $T_{\alpha}^{\beta}$ is the energy-momentum
tensor of the mathematical fluid. The definitions (\ref{rhodef})
and (\ref{pdef}) reproduce (\ref{rhoeval}) and (\ref{peval}) (now
recognized as the Friedmann equations \cite{friedmann}) from the
field equations
\begin{equation}
G_{\alpha}^{\beta}+\mathcal{C} \delta_{\alpha}^{\beta}=8 \pi
T_{\alpha}^{\beta}. \label{einsteingeneral}
\end{equation}
The question of interest here is the treatment of the constants
$\lambda_{c}$ and $\lambda_{v}$ in the procedure of applying the
energy conditions. The three distinct ways of proceeding are:

\bigskip
(i) $\mathcal{C}=\lambda_{c}-\lambda_{v} \equiv \Lambda$: First
note that $\Lambda$ is an observable (let us call it the observed
cosmological constant). The apparent smallness of $\Lambda$ arises
out of the almost perfect cancellation of $\lambda_{c}$ and
$\lambda_{v}$ an explanation of which constitutes a standard form
of one of the ``cosmological constant problems" \cite{weinberg}.
The procedure of setting $\mathcal{C}=\lambda_{c}-\lambda_{v}
\equiv \Lambda$ is the common one when energy conditions are not
under consideration and, for example, is the procedure used
(implicitly) to construct the observers
$\Omega_{\Lambda}-\Omega_{M}$ plane \cite{lake1}. As regards the
application of energy conditions, this procedure is \textit{ad
hoc} since the vacuum contribution is, without explanation,
extracted from the energy-momentum tensor and considered part of
the geometry. We write the associated energy density and isotropic
pressure of the perfect fluid source as $\rho$ and $p$
respectively. These are given by (\ref{rhoeval}) and
(\ref{peval}). (We view $\mathcal{C}=-\lambda_{v}$ as a special
case where one simply assumes, without explanation, that
$\lambda_{c}=0$.)

\bigskip
(ii) $\mathcal{C}=\lambda_{c}$: In this procedure the vacuum
contribution is not extracted from the energy-momentum tensor but
rather considered part of it. The constant $\lambda_{c}$ is not
directly observable. Rather, its value can only be inferred from
the fact that $\lambda_{c} \sim \lambda_{v}$ and with
$\lambda_{v}$ calculated as in the standard model. The perfect
fluid source now has an associated energy density and isotropic
pressure given by $\tilde{\rho} \equiv \rho-\frac{\lambda_{v}}{8
\pi}$ and $\tilde{p} \equiv p+\frac{\lambda_{v}}{8 \pi}$
respectively. These are given by (\ref{rhoeval}) and (\ref{peval})
with $\rho$ and $p$ replaced by $\tilde{\rho}$ and $\tilde{p}$.
This procedure, as regards the application of energy conditions,
is clearly more natural than (i).

\bigskip
(iii) $\mathcal{C}=0$: In this procedure the geometrical
contribution from $\lambda_{c}$ and the contribution from the
vacuum $\lambda_{v}$ are both considered part of the
energy-momentum tensor.  Exactly why the geometrical contribution
$\lambda_{c}$ should be considered part of energy-momentum tensor
is left unexplained. The perfect fluid source now has an
associated energy density and isotropic pressure given by
$\bar{\rho} \equiv \rho-\frac{\lambda_{v}-\lambda_{c}}{8 \pi}$ and
$\bar{p} \equiv p+\frac{\lambda_{v}-\lambda_{c}}{8 \pi}$
respectively. Again these are given by (\ref{rhoeval}) and
(\ref{peval}) with $\rho$ and $p$ replaced by $\bar{\rho}$ and
$\bar{p}$. (This procedure can also be considered as a special
case of (ii) where one simply assumes, again without explanation,
that $\lambda_{c}=0$.)

\bigskip
\textit{Energy Conditions.} We now impose standard energy
conditions on the mathematical fluid \cite{he} \cite{visser}
\cite{poisson}. The local energy conditions considered here are
the null energy condition (NEC), weak energy condition (WEC),
strong energy condition (SEC) and dominant energy condition (DEC)
\cite{clarification}. We let EC designate the simultaneous
satisfaction of all four energy conditions. In the present context
we are always dealing with a perfect fluid and in that case the
conditions are given in case (i) by
\begin{equation}
\hbox{NEC} \iff \quad (\rho + p \geq 0 ),
\end{equation}
\begin{equation}
\hbox{WEC} \iff \quad (\rho \geq 0 ) \hbox{ and } (\rho + p \geq
0),
\end{equation}
\begin{equation}
\hbox{SEC} \iff \quad (\rho + 3 p \geq 0 ) \hbox{ and } (\rho + p
\geq 0),
\end{equation}
and
\begin{equation}
\hbox{DEC} \iff \quad (\rho \geq 0 ) \hbox{ and } (\rho \pm p \geq
0),
\end{equation}
where $\rho$ and $p$ are replaced by $\tilde{\rho}$ and
$\tilde{p}$ in case (ii) and by $\bar{\rho}$ and $\bar{p}$ in case
(iii). From (\ref{rhoeval}) and (\ref{peval}) for case (i) we find
\begin{equation}
\rho+3p \geq 0 \iff \mathcal{C} \geq \frac{3 \ddot{a}}{a},
\end{equation}
\begin{equation}
\rho+p \geq 0 \iff  k+\dot{a}^2 \geq a \ddot{a}, \label{rhop}
\end{equation}
\begin{equation}
\rho-p \geq 0 \iff \mathcal{C} \leq \frac{2}{a^2}(k +
\dot{a}^2)+\frac{\ddot{a}}{a},
\end{equation}
and
\begin{equation}
\rho \geq 0 \iff \mathcal{C}\leq \frac{3}{a^2}(k + \dot{a}^2),
\end{equation}
again with $\rho$ and $p$ are replaced by $\tilde{\rho}$ and
$\tilde{p}$ in case (ii) and by $\bar{\rho}$ and $\bar{p}$ in case
(iii). Case (iii) has been considered in detail previously
\cite{visser1}. As we show below, this case is not of interest for
our present considerations. From the foregoing conditions it
follows that \cite{note1}
\begin{equation}
\frac{3 \ddot{a}}{a}\leq \mathcal{C} \leq \frac{2}{a^2}(k +
\dot{a}^2)+\frac{\ddot{a}}{a} \Rightarrow \hbox{EC}
\label{sufficient}.
\end{equation}

\bigskip
\textit{A bounce.} We characterize regular minima via the
conditions
\begin{equation}
\dot{a}=0,\;\;\;\ddot{a}>0,\;\;\;a>0 \label{regularminima}
\end{equation}
at, say, $t=t_{m}$. It follows immediately from (\ref{rhop}) that
$k=1$ at a regular minimum (and in particular, $k=0$ is not a
possibility). Whereas the timelike convergence condition
($R^{\alpha}_{\beta}W_{\alpha}W^{\beta} \geq 0 \;\; \forall$
timelike $W^{\alpha}$) fails at all regular minima (e.g.
$W^{\alpha}=u^{\alpha}$), it follows from (\ref{sufficient}) that
there remains in general a non-vanishing interval in $\mathcal{C}$
sufficient for the satisfaction of the energy conditions in cases
(i) with $\lambda_{c}>\lambda_{v}$ and (ii) with $\lambda_{c}>0$.
Indeed with $k=1$ and (\ref{regularminima}) it follows that the
range in $\mathcal{C}$ can be narrowed to give
\begin{equation}
\frac{3 \ddot{a}}{a}\leq \mathcal{C}\leq \frac{2}{a^2} \Rightarrow
\hbox{EC}. \label{sufficient1}
\end{equation}
The marginal case $\dot{a}=\ddot{a}=0,\;\;a>0$ distinguishes
flatness at $t=t_{m}$ ($k=\mathcal{C}=0$) and $k=1$ with the
obvious substitutions in (\ref{sufficient}) and
(\ref{sufficient1}). In both these cases the timelike convergence
condition holds \cite{maxima}.

\bigskip
\textit{Discussion.} We have shown, in a model independent way,
that in a spatially closed FLRW universe there exits a range in
$\mathcal{C}>0$ in cases (i) and (ii) that serves as a sufficient
condition for the satisfaction of the standard energy conditions
at a bounce. Since we have used no properties of the scale factor
$a$, other than the existence of a bounce, nothing can be said
about the lower limit
\begin{equation}
\mathcal{C} \geq \frac{3 \ddot{a}}{a}|_{_{t_{m}}}>0.
\label{lowerlimit}
\end{equation}
Expressing the upper limits in terms of $a(t_{m})$ we have
\begin{equation}
a(t_{m}) \leq \sqrt{\frac{2}{\Lambda}},\;\;\;\; a(t_{m}) \leq
\sqrt{\frac{2}{\lambda_{c}}} \label{upperlimit}
\end{equation}
in cases (i) and (ii) respectively. In case (i), since $\Lambda <
\; \sim 10^{-56} cm^{-2}$ the lower bound on the upper limit to
$a(t_{m})$ is of the order of $4.5\;\; 10^3$ Mpc essentially
equivalent to the current Hubble scale ($\frac{c}{H_{0}}$) and so
one might hope that the existence or not of a bounce might be
answerable without recourse to quantum cosmology. The fact that
the energy conditions need not be violated at a bounce for
cosmologies consistent with the Hubble Space Telescope
observations is certainly of interest. However, as pointed out
above, this procedure, as regards the application of energy
conditions, is \textit{ad hoc} in that it extracts $\lambda_{v}$
from the energy-momentum tensor without any explanation.

In case (ii) we must \textit{assume} that $\lambda_{c}>0$ for the
energy conditions not to be violated at a bounce. Moreover, in
contrast to case (i), if we trust the standard model all the way
to the Planck scale, then the lower bound on the upper limit to
$a(t_{m})$ is within an order or so of the Planck scale itself and
the existence or not of a bounce becomes a matter of quantum
cosmology. This procedure has to be considered the preferred one
since the role of $\lambda_{c}$ and of $\lambda_{v}$ is unaltered
by \textit{ad hoc} assumptions. Whereas the classical energy
conditions may not play a fundamental role in quantum cosmology,
the fact that they need not be violated at a bounce is certainly
of interest. Moreover, the the avoidance of energy condition
violations at a bounce is tied to the mysterious cosmological
constant problem wherein the purely classical statement
(\ref{lovelock}) is, by the observation that $\lambda{_c} \sim
\lambda_{v}$, intimately connected to the non-classical
calculation of $\lambda_{v}$.

\begin{acknowledgments}
This work was supported by a grant to KL from the Natural Sciences
and Engineering Research Council of Canada.
\end{acknowledgments}

\end{document}